A jar contains 5 red marbles numbered 1 to 5 and 8 blue marbles numbered 1 to 8 . A marble is drawn at random from the jar. Find the probability the marble is

a. Even-numbered given that the marble is red.

b. Red given that the marble is even-numbered.


1
Expert's answer
2021-11-08T17:41:21-0500

a.


"Red=\{1,2,3,4,5\}"


"Blue=\{1,2,3,4,5,6,7,8\}"


"P(even|red)=\dfrac{2}{5}"

b.


"Red=\{1,2,3,4,5\}"


"Blue=\{1,2,3,4,5,6,7,8\}"


"P(red|even)=\dfrac{2}{6}=\dfrac{1}{3}"